Find f(2) given f(x)= -4x^2+3x=29
Found 2 solutions by Cintchr, sdmmadam@yahoo.com:
Answer by Cintchr(475)

As your question is stated ... it has two = signs. I am going to assume that the =29 should actually be +29.
f(x) is just a fancy way of saying y. When given a number such as f(2), it simply means that wherever that particular variable is, substitute in the given number.
f(2)
f(x) = -4x^2 + 3x + 29
f(2) = -4(2)^2 + 3(2) + 29
f(2) = -4(4) + 6 + 29
f(2) = -16 + 6 + 29
f(2) = -10 + 29
f(2) = 19

Answer by sdmmadam@yahoo.com(530)

The problem might have been
Find f(2) given f(x)= -4x^2+3x+29 (why such a guess. The plus sign is got when shift and the equal to sign tabs are pressed and probably the shift press was not registered, because the problem cannot have two = signs
f(x)= -4x^2+3x+29----(1)
Putting x = 2 in this
f(2) = -4X(2^2)+3X(2)+29
=-4X4+6+29
=-16+6+29
=-10+29
= 19